OpenStudy (hannahc234):

Check my work If an object is dropped from a height of 144 feet, the function h(t)= -16t^2+144 gives the height of the object after t seconds. When will the object hit the ground? A. 9 s B. 6 s C. 1.5 s D. 3 s

OpenStudy (hannahc234):

I think it's D

OpenStudy (hannahc234):

@Photon336

OpenStudy (photon336):

here's why. you know that to find the time when the object hits the ground the vertical distance is going to be zero. so we can set the first equation equal to zero. \[ -16t^{2}+144 = 0 \] \[\frac{ -144 }{ -16 } = t^{2}\] \[\sqrt{\frac{ 144 }{ 16 }} = \sqrt{9} = 3 = t \]

OpenStudy (hannahc234):

Thank you